We use the t distribution to construct a confidence interval for the population mean when the underlying population standard deviation is not known. Under the assumption that the population is normally distributed, find tα/2,df for the following scenarios. (You may find it useful to reference the t table. Round your answers to 3 decimal places.)

a. A 90% confidence level and a sample of 13 observations.
b. A 95% confidence level and a sample of 13 observations.
c. A 90% confidence level and a sample of 22 observations.
d. A 95% confidence level and a sample of 22 observations.
Answer #1

Solution :

Using t table,

(a)

sample size = n = 13

Degrees of freedom = df = n - 1 = 12

t /2,df = t0.05,12 = 1.782

(b)

sample size = n = 13

Degrees of freedom = df = n - 1 = 12

t /2,df = t0.025,12 = 2.179

(c)

sample size = n = 22

Degrees of freedom = df = n - 1 = 21

t /2,df = t0.05,21 = 1.721

(d)

sample size = n = 22

Degrees of freedom = df = n - 1 = 21

t /2,df = t0.025,21 = 2.080